OpenStudy (anonymous):

will give metal and become come fan to best answer Multiply (2.1 x 10^8) (4 x 10^5) express the result in scientific notation A. 8.4 x 10^12 B. 8.4 x 10^13 C. 8.4 x 10^39 D. 8.4 x 10^40

OpenStudy (dan815):

what kind of metal?

OpenStudy (anonymous):

? what do you mean?

OpenStudy (dan815):

what's a come fan

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

if you are not going to answer the question then please don't reply

OpenStudy (dan815):

10^2 = 10 * 10

OpenStudy (dan815):

what is 10*10

OpenStudy (anonymous):

100

OpenStudy (dan815):

so 10^2 = 100

OpenStudy (dan815):

what is 10^3?

OpenStudy (anonymous):

1000

OpenStudy (dan815):

good

OpenStudy (dan815):

what is 10 * 10 = 10^1 * 10^1 = 10^(1+1) = 10^2

OpenStudy (anonymous):

i have not a clue.....

OpenStudy (dan815):